standfast01:

Rachel must store a few pieces of her furniture for six months. A storage unit costs $55 for the first month and $47.50 for the remaining months. How much will it cost to store her furniture? A. $237.50 B. $285.00 C. $292.50 D. $330.00

JustSaiyan:

Add $55 to $47.50 · 5 $47.50 · 5 = 237.50 + 55 = $292.50

standfast01:

Thx

JustSaiyan:

You're welcome.